Ultra Ltd is a chip manufacturer based in Taiwan. Its stock traded at $42 at the beginning of 2018 with a beta estimate of 1. Its market risk premium is 6%. The risk free rate at the end of 2017 was 1.6%. The company is expected to pay dividends of $1.60 per share at the end of 2018 and 2019.

(i) Using the Capital Asset Pricing Model (CAPM), calculate the required rate of return.

(ii) What will be the Ultra Ltd’s share price at the end of 2019 if you forecast that the company will not pay the dividends as expected?

(iii) What will be Ultra Ltd’s share price at the end of 2019 if you forecast that the company will pay the dividends as expected?

(iv) Assuming that Ultra Ltd’s investors expect it to pay a dividend of $2.50 per share forever. Using the required rate of return calculated in (i) above, what would be the value gained or lost per share if you had bought the shares at $42.

(v) Ultra Ltd traded at 4.5 times sales in 2017 and it reported a net profit margin of its sales of 14%. Determine its P/E ratio at the end of 2017.

(vi) Ultra Ltd had 700,000 shares at the end of 2017. The company issued an additional 250,000 shares to the market at the market price of $42 per share on 20 January 2018. What will be the effect of this share issue on the price per share.

(vii) The directors of Ultra Ltd decided to issue some new preferred stock on 7 April 2018. The issue will pay a $20 annual dividend per share, beginning 20 years from the date of issuance. If the market requires a 6% return on this investment, how should the preferred stock be priced on issuance?
Answer #1

Share price = $ 42 beginning 2018

Beta estimate= 1

Market risk premium or (Rm - Rf) = 6%

Risk free rate = 1.6%

Expected dividend= $ 1.6 at year end 2018

A) As per CAPM

Required rate of return = Rf + (Rm - Rf) / Beta

= 1.6 + 6/1

=7.6%

B) Price at year 2019 end if dividends are not paid as expected

Price at beginning of 2018 = 42

End of 2019 price =price / (1 + Ke)2 Ke= Cost of Equity or required rate of return

= 42 /(1 + 7.6%) = 42 / 1.076)2 = 36.276

C) Price at year 2019 end if dividends are paid as expected

Price at beginning of 2018 = 42

End of 2019 price =price + expected dividend / (1 + Ke)2 Ke= Cost of Equity or required rate of return

= 42 + 1.6 /(1 + 7.6%) = 43.6 / 1.076)2 = 37.658

D) The value of stock (if dividend is paid forever) = dividend / Ke

= 2.5 / 7.6%

=2.5/ 0.076= 32.89

Value lost per share = 42 - 32.89 = 9.11

E) Ultra Ltd trading at 4.5 times of its sales

sales = 42/4.5 = 9.33

Earnings = 14 %

Earnings = 9.33 * 14 % = 1.306

P/E ratio = Price / Earnings

= 42/1.306 =32.15